# Universality conditions of unified classical and quantum reservoir computing

### v2

@@ Abstract
Reservoir computing is a versatile paradigm in computational neuroscience and machine learning, that exploits a recurrent neural network to efficiently process time-dependent information. The power of many neural network architectures resides in their universality approximation property. As widely known, classes of reservoir computers serve as universal approximators of functionals with fading memory. The construction of such universal classes often appears context-specific, but in fact, they follow the same principles. Here we present a unified theoretical framework and we propose a ready-made setting to secure universality, based on the minimal sufficient conditions for a class of reservoir computers to be universal, namely the fading memory and the polynomial algebra structure of the set of their associated functionals. We test the result in the arising context of quantum reservoir computing. Guided by such a unified theorem we suggest why spatial multiplexing serves as a computational resource when dealing with quantum registers, as empirically observed in specific implementations on quantum hardware. The analysis sheds light on a unified view of classical and quantum reservoir computing.
@@ 1 Introduction
Reservoir computing is a general computational framework that exploits the nonlinear dynamics of a recurrent neural network and simple readout functions for the online processing of time-dependent inputs (Nakajima et al. (2019); Jaeger et al. (2007)). The spread of reservoir computing has stemmed from two primary paradigms, namely echo state networks (ESN), for time series prediction, (Jaeger (2001); Jaeger and Haas (2004)) and liquid state machines (LSM) respectively, the latter inspired by spiking neuronal activity for modeling online time-dependent inputs (Maass et al. (2002, 2007)).
@@
A computational architecture based on neural networks is said to be universal if for any given target function there exists an instance of the architecture realizing a mapping between the input and the output that approximates the given function with arbitrary precision. The issue of finding neural networks that serve as universal approximators has attracted considerable mathematical research during the years (Kolmogorov (1956); Boyd and Chua (1985); Sandberg (1991); Matthews (1993); Cucker and Smale (2002); Cucker and Zhou (2007)), paving the way for the implementation of versatile machine learning paradigms. Universal approximation results have been proven for different neural network architectures, starting from the classical results for feed-forward (Hornik et al. (1989); Cybenko (1989); Hornik (1991, 1993); Sprecher (1996, 1997); Stinchcombe (1999)) and for recurrent neural networks, respectively (Hammer (2000); Schäfer and Zimmermann (2006)). More recently, universality has been proven also for invertible (Jin et al. (2024); Ishikawa et al. (2024)), for mean-field (Pham and Warin (2023)), and for deep convolutional neural networks (Zhou (2020)).
@@
Restricting to reservoir computing, both ESN (Grigoryeva and Ortega (2018b, a); Gonon and Ortega (2020)) and LSM (Maass and Markram (2004) have yielded classes of reservoir computers that serve as universal approximators of functionals with fading memory. Despite the similarity in the strategy exploited in the proofs of these results, ultimately based on an application of the Stone-Weierstrass theorem in the spirit of the seminal work of Boyd and Chua in 1985 (Boyd and Chua (1985)), they appeared as context-specific.
@@
Emerging from these frameworks, as anticipated by Maass (Maass (2010)), reservoir computing has proved its efficiency not restricted to a digital approach. In the last twenty years, numerous physical implementations of reservoir systems have been proposed as valid tools for reservoir computing (Lukoševičius and Jaeger (2009); Tanaka et al. (2019); Marković et al. (2020); Nakajima (2020); Lee et al. (2023)), exploiting for example, optical devices (Larger et al. (2012); Vandoorne et al. (2014); Brunner et al. (2021); Nakajima et al. (2021)), spiking neurons and cortical networks (Dockendorf et al. (2009); Yada et al. (2021); Venker et al. (2023); Peng et al. (2024)) and nanoscale oscillators (Torrejon et al. (2017); Papp et al. (2021)). More recently, the advent of quantum computation (Prati (2017); De Michielis et al. (2023)) has also shed light on the development of reservoir computers harnessing the high computational capacity of quantum reservoir systems. After the first proposal (Fujii and Nakajima (2017)), many quantum systems have been proposed as reliable effective reservoirs (Kutvonen et al. (2020); Govia et al. (2021); Martínez-Peña et al. (2021); Mujal et al. (2021); Suzuki et al. (2022); Martínez-Peña and Ortega (2022); Mujal et al. (2023); Molteni et al. (2023); Ghosh et al. (2021b); García-Beni et al. (2023); Domingo et al. (2023)) and even as examples of universal classes of reservoir computers (Chen and Nurdin (2019); Chen et al. (2020); Nokkala et al. (2021)). In particular, quantum networks have been proven effective in processing physical quantum information and predicting quantum dynamics (Ghosh et al. (2019, 2021b, 2021a); Lazzarin et al. (2022)). As new embodiments of reservoir computing are emerging after the advent of quantum computing, a unified framework is needed in order to assign universality. This work aims to define such a generalized framework to determine which minimum set of properties is mandatory to be fulfilled by a family of reservoir computers to exhibit universality. A unified theoretical approach, that includes systems fed with inputs living in general normed space, may act as a guideline promoting the development of new implementations of reservoir systems.
@@
Due to the large variety of reliable classes of reservoir computers, we propose a unified theoretical framework - that we call echo reservoir computers - together with a ready-made setting that guarantees universality. We then show that echo state networks and liquid state machines fall within our unified framework and we discuss an example that entails a quantum reservoir (Nakajima et al. (2019)).
@@
It results that the sufficient conditions for a class of reservoir computers to be universal are respectively that the set of associated functionals has continuity with respect to a fading metric and is a polynomial algebra. The fading metric accounts for the fading memory property by combining the $l^{\infty}$ norm with a monotonic increasing function in $\mathbb{R}^{-}$, referred to as the fading function. In this perspective, as a study case, we test such a unified framework considering spatial multiplexing, namely by parallelly processing information with disjoint reservoirs and inter-depending readout functions. Already empirically (Nakajima et al. (2019)) and theoretically (Chen and Nurdin (2019); Chen et al. (2020); Nokkala et al. (2021) recognized as a relevant booster for the computational power of quantum reservoir systems, spatial multiplexing appears as a resource for building universal quantum reservoir systems also within our general framework.
@@ 2.1 Definitions of the input data and the reservoir
As a consequence of the latter lemma, one is allowed to consider only real functionals, which are simpler mathematical objects with respect to filters. In particular, it allows the application of Stone - Weierstrass theorem for filters and functionals with target in $\mathbb{R}$.
@@ 2.2 Universal echo reservoir computers
Let $\mathcal{N}_{1},\mathcal{N}_{2}$ be two normed spaces. Let $E\colon\mathcal{N}_{1}^{T}\rightarrow\mathcal{N}_{2}^{T}$ be a given BCTI filter and $H\colon\mathcal{N}_{2}\rightarrow\mathbb{R}$ a given real functional. We call echo reservoir computer the system defined by
@@
Denote with $T_{-}$ either $\mathbb{Z}_{-}$ or $\mathbb{R}_{-}$. We call restricted orbit any function of the form $u\colon T_{-}\rightarrow\mathbb{R}^{n}$. Accordingly to Def. 2.2, we denote with $S^{T,\mathcal{N}}_{K,-}$ the set of all the bounded orbits and with $I_{-}\subset S^{T,\mathcal{N}}_{K,-}$ any subset.
@@
Note that the unique functional associated with a BCTI filter is uniquely determined by its action on restricted orbits.
@@
A family of functionals $\mathcal{H}$ separates orbits if for any orbits $u,v\in I_{-}$ with $u\not=v$ there exists a functional $H\in\mathcal{H}$ such that $Hu\not=Hv$.
@@
    any functional $R\in\mathcal{H}$ is continuous with respect to the metric $d$.
@@
In the latter theorem, we ask the family of functionals associated with a given class of reservoir computers to be a polynomial algebra. In many examples, one exploits polynomial readouts to ensure this condition is verified, although is not necessary. See Refs. Grigoryeva and Ortega (2018b, a) for examples of universal classes of reservoir computers with linear readout.
@@ 2.3 Continuity and fading memory
In the context of recurrent neural networks, the concept of fading memory plays a key role (Boyd and Chua (1985); Sandberg (2003)). Informally, a functional, or a filter, has fading memory if its dependence on older information is reduced over time. In the former subsection, we considered the sufficient conditions for the approximation of continuous functionals defined on inputs belonging to compact metric space. In this section, we show that fading memory provides suitable continuity to be assured by filters and functionals. We specify its formal definition for two different sets of input orbits, namely time series and spike train.
@@
We consider first time series, thus bounded orbits labeled in $T=\mathbb{Z}$, namely $I_{-}^{K}=S^{\mathbb{Z},\mathcal{N}}_{K}$. They lay in a compact metric space for a proper "fading" distance.
@@
$\leq\sup_{t\in T_{-}}\left(\norm{u_{t}-v_{t}}+{v_{t}-z_{t}}\right)\omega_{t}\leq$
@@
We prove that $\left(I_{-}^{K},d_{\omega}\right)$ is compact by proving that one can extract a convergent subsequence from any sequence in it. For convenience, for any $m\in\mathbb{N}$ and any $u=(u_{-\infty},\dots,u_{-1},u_{0})\in I_{-}^{K}$ we denote with $u^{m}=(u_{-m},\dots,u_{-1},u_{0})$ the time-restricted orbit and with $I_{-,m}^{K}$ the space of these time-restricted orbits. Then, any $I_{-,m}^{K}$ is compact, and thus sequentially compact, since it is bounded and closed in the finite-dimensional normed space $\mathcal{N}^{m}$.
Given $\left\{u(n)\right\}_{n\in\mathbb{N}}$ a sequence in $I_{-}^{K}$, we extract from it a subsequence $u(n_{j})$ such that $d_{\omega}(u(n_{j}),\tilde{u})\rightarrow 0$ for a certain $\tilde{u}\in I_{-}^{K}$. For any fixed $k\in\mathbb{N}$, $u^{k}(n)$ is a sequence in $I_{-,m}^{K}$. Thus, since $I_{-,m}^{K}$ is sequentially compact, there exists a sequence $n_{k}$ such that
$\sup_{t\in\{-m,\dots,0\}}\norm{u^{k}(n_{k})-\tilde{u}^{k}}\rightarrow 0,\quad n_{k}\rightarrow+\infty\,.$
@@
We come to prove that $d_{\omega}\left(u(n_{j}),\tilde{u}\right)\rightarrow 0$ as $j\rightarrow+\infty$. Fix $\epsilon>0$. By definition of fading function $\omega$, there exists $T$ such that for any $t\geq T$
$\sup_{t\leq-T}\norm{u_{t}(n_{j})-\tilde{u}_{t}}\omega_{-t}\leq 2K\omega_{T}\leq\epsilon$ (3)
@@
Given any fading function $\omega$, one can construct a distance $d^{c}_{\omega}$ for which fading memory functionals defined on $(I^{\Delta},d^{c}_{\omega})$ are continuous (Maass and Markram (2004)). To do this, we approximate any $u\in I^{\Delta}$ with a continuous function $f^{u}$. Explicitly, one can define such a function by
@@
The following result is proved in Ref. Maass and Markram (2004).
@@ 3 Universality of classical reservoir computers
The definitions and the theorem above are exploited in the following, as they provide the minimal key ingredients for constructing classes of universal reservoir computers. First, the two classical paradigms of reservoir computing, namely liquid state machine and echo state networks (Maass and Markram (2004); Grigoryeva and Ortega (2018b)) are considered, and next the unified paradigm is applied to an example of a quantum reservoir (Nakajima et al. (2019)).
@@ 3.1 Universality of classical echo state networks
In Ref. Grigoryeva and Ortega (2018a) it is identified a ready-to-use condition that assures the echo state property of an ESN and the fading memory of the associated functional. We report this lemma for completeness since we use it later on in our developments.
@@ 3.2 Recasting universality of liquid state machines in the unified framework
with $x_{t}\in\mathbb{R}^{n}$ and $B=(B_{1},\dots,B_{N})$ a N-dimensional causal and time-invariant filter. If the input functions $u$ are continuous, uniformly bounded, and uniformly Lipschitz functions of time, many implementations of the family of filters $B$ and readout $h$ are possible, so that separation of points, constants representation and fading memory are ensured. Then, the corresponding class of liquid state machines is universal, as a consequence of Theorem 2.15. We refer to Maass and Markram (2004) for a precise statement.
Here we discuss an implementation of an LSM in the spirit of neural microcircuit networks, considering a reservoir composed of a network of $N$ connected integrate-and-fire spiking neurons (Maass et al. (2002)). The current liquid state $x_{t}$ is an N-dimensional vector that represents the contribution of each neuron to the membrane potential at time $t$. Formally, we can describe the associated filters as linear operators with exponential decay (see Maass et al. (2002) for more details) taking spike train inputs $u\in I^{\Delta}$
@@
where we are denoting with $\delta_{u}$ the delta function $\delta_{u}(t)\coloneqq\sum_{s\in u}\delta(t-s)$, and with $b_{i}(s)$ a continuous, bounded, real function. Such a filter has fading memory with $\omega_{t}=e^{t}$. Moreover, it is causal and time-invariant by construction. The readout function $h$ is implemented through a subset of firing neurons that receive connections from any neuron in the reservoir, but that are not mutually connected. Concretely, the current firing activity of this subset of neurons is interpreted as the analogical output of the readout function. It is known that such functions $h$ are universal approximators, namely, they approximate with arbitrary precision any given real function (Maass (2000)).

### v3

@@ Abstract
Reservoir computing is a versatile paradigm in computational neuroscience and machine learning, that exploits a recurrent neural network to efficiently process time-dependent information. The power of many neural network architectures resides in their universality approximation property. As widely known, classes of reservoir computers serve as universal approximators of functionals with fading memory. The construction of such universal classes often appears context-specific, but, in fact, they follow the same principles. Here we present a unified theoretical framework and we propose a ready-made setting to secure universality, based on the minimal sufficient conditions for a class of reservoir computers to be universal, namely the fading memory and the polynomial algebra structure of the set of their associated functionals. We test the result in the arising context of quantum reservoir computing. Guided by such a unified theorem we suggest why spatial multiplexing serves as a computational resource when dealing with quantum registers, as empirically observed in specific implementations on quantum hardware. The analysis sheds light on a unified view of classical and quantum reservoir computing.
@@ 1 Introduction
Reservoir computing is a general computational framework that exploits the nonlinear dynamics of a recurrent neural network and simple readout functions for the online processing of time-dependent inputs [1, 2]. The spread of reservoir computing has stemmed from two primary paradigms, namely echo state networks (ESN), for time series prediction, [3, 4] and liquid state machines (LSM) respectively, the latter inspired by spiking neuronal activity for modeling online time-dependent inputs [5, 6].
@@
A computational architecture based on neural networks is said to be universal if for any given target function there exists an instance of the architecture realizing a mapping between the input and the output that approximates the given function with arbitrary precision. The issue of finding neural networks that serve as universal approximators has attracted considerable mathematical research during the years [7, 8, 9, 10, 11, 12], paving the way for the implementation of versatile machine learning paradigms. Universal approximation results have been proven for different neural network architectures, starting from the classical results for feed-forward [13, 14, 15, 16, 17, 18, 19] and for recurrent neural networks, respectively [20, 21]. More recently, universality has been proven also for invertible [22, 23], for mean-field [24], and for deep convolutional neural networks [25].
@@
Restricting to reservoir computing, both ESN [26, 27, 28] and LSM [29] have yielded classes of reservoir computers that serve as universal approximators of functionals with fading memory. Despite the similarity in the strategy exploited in the proofs of these results, ultimately based on an application of the Stone-Weierstrass theorem in the spirit of the seminal work of Boyd and Chua in 1985 [8], they appeared as context-specific.
@@
Emerging from these frameworks, as anticipated by Maass [30], reservoir computing has proved its efficiency not restricted to a digital approach. In the last twenty years, numerous physical implementations of reservoir systems have been proposed as valid tools for reservoir computing [31, 32, 33, 34, 35], exploiting for example, optical devices [36, 37, 38, 39], spiking neurons and cortical networks [40, 41, 42, 43] and nanoscale oscillators [44, 45]. More recently, the advent of quantum computation [46, 47] has also shed light on the development of reservoir computers harnessing the high computational capacity of quantum reservoir systems. After the first proposal [48], many quantum systems have been proposed as reliable effective reservoirs [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] and even as examples of universal classes of reservoir computers [60, 61, 62]. In particular, quantum networks have been proven effective in processing physical quantum information and predicting quantum dynamics [63, 57, 64, 65]. As new embodiments of reservoir computing are emerging after the advent of quantum computing, a unified framework is needed in order to assign universality. This work aims to define such a generalized framework to determine which minimum set of properties is mandatory to be fulfilled by a family of reservoir computers to exhibit universality. A unified theoretical approach, that includes systems fed with inputs living in general normed space, may act as a guideline promoting the development of new implementations of reservoir systems.
@@
Due to the large variety of reliable classes of reservoir computers, we propose a unified theoretical framework - that we call echo reservoir computers - together with a ready-made setting that guarantees universality. We then show that echo state networks and liquid state machines fall within our unified framework and we discuss an example that entails a quantum reservoir [1].
@@
It results that the sufficient conditions for a class of reservoir computers to be universal are respectively that the set of associated functionals has continuity with respect to a fading metric and is a polynomial algebra. The fading metric accounts for the fading memory property by combining the sup-norm with a monotonic increasing function in $\mathbb{R}^{-}$, referred to as the fading function. In this perspective, we test such a unified framework, showing that it encompasses the two most notable embodiments of reservoir computers (ESN and LSM). Moreover, we prove that our framework is general, including an example concerning a gate-based quantum reservoir. In this context, spatial multiplexing, namely the parallel processing of information with disjoint registers and inter-depending readout functions, accounts for universality, as empirically recognized in [1] and theoretically confirmed in [60, 61, 62].
@@ 2.1 Definitions of the input data and the reservoir
As a consequence of the latter lemma, one is allowed to consider only functionals, which are simpler mathematical objects with respect to filters. In particular, it allows the application of Stone - Weierstrass theorem for filters and functionals with target in $\mathbb{R}$.
@@ 2.2 Universal echo reservoir computers
Let $\mathcal{N}_{1},\mathcal{N}_{2}$ be two normed spaces. Let $E\colon\mathcal{N}_{1}^{T}\rightarrow\mathcal{N}_{2}^{T}$ be a given BCTI filter and $H\colon\mathcal{N}_{2}\rightarrow\mathbb{R}$ a given real function. We call echo reservoir computer the system defined by
@@
Denote with $T_{-}$ either $\mathbb{Z}_{-}$ or $\mathbb{R}_{-}$. We call restricted orbit any function of the form $u\colon T_{-}\rightarrow\mathbb{R}^{n}$. Accordingly to Def. 2.2, we denote with $S^{T,\mathcal{N}}_{K,-}$ the set of all the bounded orbits and with $I_{-}\subset S^{T,\mathcal{N}}_{K,-}$ any subset of inputs.
@@
The unique functional associated with a BCTI filter is uniquely determined by its action on restricted orbits.
@@
A family of functionals $\mathcal{H}$ separates orbits in $I_{-}$ if for any pair of orbits $u,v\in I_{-}$ with $u\not=v$ there exists a functional $H\in\mathcal{H}$ such that $Hu\not=Hv$.
@@
    any functional $R\in\mathcal{H}$ is continuous with respect to the metric $d$
@@
In the latter theorem, we ask the family of functionals associated with a given class of reservoir computers to be a polynomial algebra. In many examples, one exploits polynomial readouts to ensure this condition is verified, although is not necessary. See Refs. [26, 27] for examples of universal classes of reservoir computers with linear readout.
@@ 2.3 Continuity and fading memory
In the context of recurrent neural networks, the concept of fading memory plays a key role [8, 66]. Informally, a functional, or a filter, has fading memory if its dependence on older information is reduced over time. In the former subsection, we considered the sufficient conditions for the approximation of continuous functionals defined on inputs belonging to compact metric space. In this section, we show that fading memory provides suitable continuity to be assured by filters and functionals. We specify its formal definition for two different sets of input orbits, namely time series and spike train.
@@
We consider first time series, thus bounded orbits labeled in $T=\mathbb{Z}$, namely $I_{-}^{K}=S^{\mathbb{Z},\mathcal{N}}_{K}$. Fading memory ensures the compactness of the associated metric space.
@@
$\leq\sup_{t\in T_{-}}\left(\norm{u_{t}-v_{t}}+\norm{v_{t}-z_{t}}\right)\omega_{t}\leq$
@@
We prove that $\left(I_{-}^{K},d_{\omega}\right)$ is compact by proving that one can extract a convergent subsequence from any sequence in it.
For convenience, for any $m\in\mathbb{N}$ and any $u=(u_{-\infty},\dots,u_{-1},u_{0})\in I_{-}^{K}$ we denote with $u^{m}=(u_{-m},\dots,u_{-1},u_{0})$ the time-restricted orbit and with $I_{-,m}^{K}$ the space of these time-restricted orbits. Then, any $I_{-,m}^{K}$ is compact, and thus sequentially compact, since it is bounded and closed in the finite-dimensional normed space $\mathcal{N}^{m}$.
Given $\left\{u(n)\right\}_{n\in\mathbb{N}}$ a sequence in $I_{-}^{K}$, we extract from it a subsequence $u(n_{j})$ such that $d_{\omega}(u(n_{j}),\tilde{u})\rightarrow 0$ for a certain $\tilde{u}\in I_{-}^{K}$. For any fixed $k\in\mathbb{N}$, $u^{k}(n)$ is a sequence in $I_{-,k}^{K}$. Thus, since $I_{-,k}^{K}$ is sequentially compact, there exists a sequence $n_{k}$ such that
$\sup_{t\in\{-k,\dots,0\}}\norm{u_{t}^{k}(n_{k})-\tilde{u}_{t}^{k}}\rightarrow 0,\quad n_{k}\rightarrow+\infty\,.$
@@
We come to prove that $d_{\omega}\left(u(n_{j}),\tilde{u}\right)\rightarrow 0$ as $j\rightarrow+\infty$. Fix $\epsilon>0$. By definition of fading function $\omega$, there exists $T_{0}$ big enough such that for any $t\geq T_{0}$
$\sup_{t\leq-T_{0}}\norm{u_{t}(n_{j})-\tilde{u}_{t}}\omega_{-t}\leq 2K\omega_{T_{0}}\leq\epsilon$ (3)
@@
Given any fading function $\omega$, one can construct a distance $d^{c}_{\omega}$ for which fading memory functionals defined on $(I^{\Delta},d^{c}_{\omega})$ are continuous [29]. To do this, we approximate any $u\in I^{\Delta}$ with a continuous function $f^{u}$. Explicitly, one can define such a function by
@@
The following result is proved in Ref. [29].
@@ 3 Universality of classical reservoir computers
The definitions and the theorem above are exploited in the following, as they provide the minimal key ingredients for constructing classes of universal reservoir computers. First, the two classical paradigms of reservoir computing, namely liquid state machine and echo state networks [29, 26] are considered, and next the unified paradigm is applied to an example of a quantum reservoir [1].
@@ 3.1 Universality of classical echo state networks
In Ref. [27] it is identified a ready-to-use condition that assures the echo state property of an ESN and the fading memory of the associated functional. We report this lemma for completeness since we use it later on in our developments.
@@
Remark 3.6.
Typically, in practical implementations, linear readouts are sufficient to perform many specific tasks. Indeed, even the class of ESN with linear readouts only (so, $p(x_{t})=Cx_{t}+D$ in (6)) has been proven to be universal [26, 27].
@@ 3.2 Recasting universality of liquid state machines in the unified framework
with $x_{t}\in\mathbb{R}^{n}$ and $B=(B_{1},\dots,B_{N})$ a N-dimensional causal and time-invariant filter. If the input functions $u$ are continuous, uniformly bounded, and uniformly Lipschitz functions of time, many implementations of the family of filters $B$ and readout $h$ are possible, so that separation of points, constants representation and fading memory are ensured. Then, the corresponding class of liquid state machines is universal, as a consequence of Theorem 2.15. We refer to [29] for a precise statement.
Here we discuss an implementation of an LSM in the spirit of neural microcircuit networks, considering a reservoir composed of a network of $N$ connected integrate-and-fire spiking neurons [5]. The current liquid state $x_{t}$ is an N-dimensional vector that represents the contribution of each neuron to the membrane potential at time $t$. Formally, we can describe the associated filters as linear operators with exponential decay (see [5] for more details) taking spike train inputs $u\in I^{\Delta}$
@@
where we are denoting with $\delta_{u}$ the delta function $\delta_{u}(t)\coloneqq\sum_{s\in u}\delta(t-s)$, and with $b_{i}(s)$ a continuous, bounded, real function. Such a filter has fading memory with $\omega_{t}=e^{t}$. Moreover, it is causal and time-invariant by construction. The readout function $h$ is implemented through a subset of firing neurons that receive connections from any neuron in the reservoir, but that are not mutually connected. Concretely, the current firing activity of this subset of neurons is interpreted as the analogical output of the readout function. It is known that such functions $h$ are universal approximators, namely, they approximate with arbitrary precision any given real function [67].